# Forgetting-MarI: LLM Unlearning via Marginal Information Regularization

### v1

@@ 4.4 Empirical verification of theoretical guarantees
Theorems 2.2 and 2.1 imply that, after Forgetting-MarI, the mutual information between logits and the “seen/unseen” bit $Z$ is negligible; hence any confidence-based test (perplexity, cross-entropy, log-likelihood, etc.) should fail to separate forgotten from genuinely unseen text.
Note: ppl = perplexity
@@
Figure 6: Detector performance for a GPT2-LG model without unlearning (left), unlearned with Forgetting-MarI (middle), and the golden-standard unlearn baseline (right).
@@ 5 Conclusion
Second, our work highlights the importance of parameter selection in unlearning effectiveness. Developing principled approaches for optimal parameter tuning, especially with theoretical guidance, remains an open challenge that could significantly enhance the practicality of unlearning methods. Additionally, future work could explore the scalability of our approach to even larger models and datasets, investigate our method’s robustness across different model architectures and domains, and the principles of our approach could be applied to models trained on other data modalities.
@@ Appendix A Appendix of Section 1
A.1 Details of LLM Unlearning Methods: Implicit Marginal Information Unlearning
@@ Appendix B Appendix of Section 2
Proof.
To start, define the Bayes error as
@@
It remains to show that the upper bound is tight. Indeed, fix an arbitrary $I\in[0,H_{2}(\pi)]$. Choose $p^{\star}\in\bigl[\tfrac{1}{2},1\bigr]$ such that $H_{2}(p^{\star})=H_{2}(\pi)-I$. Construct $P_{Z\mid X_{\operatorname{MarI}}}$ such that $P(Z=1\mid X_{\operatorname{MarI}})\in\{p^{\star},\,1-p^{\star}\}$ with probabilities chosen to match the prior $\pi$. Then $H(Z\mid X_{\operatorname{MarI}})=H_{2}(p^{\star})$ and $I(X_{\operatorname{MarI}};Z)=I$, while the Bayes error satisfies $P_{e}=\min\{p^{\star},1-p^{\star}\}=H_{2}^{-1}(H(Z)-I)$. Hence, equality holds in the bound. ∎
@@ B.2 Why Mutual Information Rather Than KL Divergence
Proof.
By the mean value theorem, for each $t$ there exists $\xi_{t}\in[\min\{p_{t}^{\,u}(u_{t}),p_{t}^{\,r}(u_{t})\},1]\subseteq[\gamma,1]$ such that
@@
Combining yields the claim. ∎
B.4 Proof of Theorem 2.2
We start with the following three lemmata that are needed for the proof of Theorem 2.2:
Lemma B.1 (Point-wise KL bound).
@@
Proof.
Fix $x\in V$ and set $y:=\dfrac{p(x)}{q(x)}\in[1,M]$. Inequality equation 4 is equivalent to
@@
Multiplying both sides by y1y-1 yields equation 4, which is precisely equation 4 after reinstating $y=p(x)/q(x)$. Therefore, equation 4 holds for every $x\in V$. This completes the proof. ∎
Lemma B.2.
@@
Proof.
Let $m=\tfrac{p+q}{2}$. Pinsker’s inequality gives $\|p-m\|_{1}^{2}\leq 2\,D_{\mathrm{KL}}(p\|m)$ and analogously for $q$. Hence
@@
Lemma B.3 (Exact TV scaling under mixture).
@@
Proof.
$p^{d}-p^{r}=(1-\alpha)(p^{u}-p^{r})$. Taking $\ell_{1}$-norms and dividing by $2$ yields the identity. ∎
Now, we are ready to prove Theorem 2.2:
Proof.
Define $Y_{t}:=\log\frac{p_{t}^{\,r}(U_{t})}{p_{t}^{\,u}(U_{t})}$, so that
@@
By Jensen’s inequality, $\frac{1}{T}\sum_{t}\sqrt{\operatorname{\mathrm{JSD}}(p_{t}^{\,d},p_{t}^{\,r})}\leq\sqrt{\frac{1}{T}\sum_{t}\operatorname{\mathrm{JSD}}(p_{t}^{\,d},p_{t}^{\,r})}.$ Combining the displays proves the claim with $I(X_{\operatorname{MarI}};Z)=\frac{1}{T}\sum_{t}\operatorname{\mathrm{JSD}}(p_{t}^{\,d},p_{t}^{\,r})$. ∎
@@ Appendix C Appendix of Section 3
C.1 Position-wise vs. Pooled MarI: Empirical Comparison
@@
A deeper theoretical and experimental analysis of these behaviors, hybrids that combine both signals, is left for future work. In our experiments, because the text is heterogeneous in both length and context, we use random mini-batches and the pooled estimator by default.
@@ Appendix D Appendix of Section 4
D.1 More Robustness Test Results
In Figure 9 we show that $10/90$ split unlearning training curve, which supplements the $50/50$ split unlearning training curve in Figure 3 of Section 4.
Figure 9: Training curves of F-MarI for the 10/90 split unlearning with the GPT2-LG model. When you do not know the true underlying stopping criteria, we choose to stop when the detector fails. This corresponds to F-Mari, lambda = 0.95 and epoch = 10.
In Figure 10 we show the training curve using a smaller learning rate than the result shown in Figure 3.
(a) Accuracy versus training epochs.
(b) Unlearning results using different regularization with stopping E.
Figure 10: 50/50 split test, F-MarI training curves using small learning rate using GPT2-LG to better observe the convergence process. Both plot shows stable curve and robustness against lambda choice. The stopping criteria is when validation accuracy drop 3 percent, with the corresponding epoch denoted as ‘E’
D.2 Supplementary General Model Capacity Test Results

### v2

@@ 4.4 Detector evaluation: Empirical verification of theoretical guarantees
Theorems 2.1 implies that, after Forgetting-MarI, the mutual information between logits and the “seen/unseen” bit $Z$ is negligible; hence any confidence-based test (perplexity, cross-entropy, log-likelihood, etc.) should fail to separate forgotten from genuinely unseen text.
@@
Figure 5: Detector performance for a GPT2-LG model without unlearning (left), unlearned with Forgetting-MarI (middle), and the golden-standard unlearn baseline (right). Here, ppl = perplexity.
@@ 5 Conclusion
Second, our work highlights the importance of parameter selection in unlearning effectiveness. Developing principled approaches for optimal parameter tuning, especially with theoretical guidance, remains an open challenge that could significantly enhance the practicality of unlearning methods. Additionally, future work could explore the scalability of our approach to even larger models and datasets, investigate our method’s robustness across different model architectures and domains, replace retain set with new data set to remove retain access assumption or for more robust fine-tuning, and the principles of our approach could be applied to models trained on other data modalities.
@@ Appendix A Appendix of Section 1
A.1 Details of LLM Unlearning Methods: Implicit Marginal Information Unlearning via conflicting forces
@@ Appendix B Appendix of Section 2
Proof To start, define the Bayes error as
@@
It remains to show that the upper bound is tight. Indeed, fix an arbitrary $I\in[0,H_{2}(\pi)]$. Choose $p^{\star}\in\bigl[\tfrac{1}{2},1\bigr]$ such that $H_{2}(p^{\star})=H_{2}(\pi)-I$. Construct $P_{Z\mid X_{\operatorname{MarI}}}$ such that $P(Z=1\mid X_{\operatorname{MarI}})\in\{p^{\star},\,1-p^{\star}\}$ with probabilities chosen to match the prior $\pi$. Then $H(Z\mid X_{\operatorname{MarI}})=H_{2}(p^{\star})$ and $I(X_{\operatorname{MarI}};Z)=I$, while the Bayes error satisfies $P_{e}=\min\{p^{\star},1-p^{\star}\}=H_{2}^{-1}(H(Z)-I)$. Hence, equality holds in the bound.
@@ B.2 Why Mutual Information Rather Than KL Divergence
Proof By the mean value theorem, for each $t$ there exists $\xi_{t}\in[\min\{p_{t}^{\,u}(u_{t}),p_{t}^{\,r}(u_{t})\},1]\subseteq[\gamma,1]$ such that
@@
Combining yields the claim.
B.4 Generalization of Theorem 2.1 to unlearn set neighborhood
Here, we show that the self-perplexity gap guarantee provided by Theorem 2.1 can be generalized to a neighborhood of $u$ rather than $u$ itself:
Theorem B.1 (MarI controls neighborhood-perplexity gap)
Draw $U:=\{U_{t}\}_{t=1}^{T}$ with $U_{t}\sim p_{t}^{\,u}$ independently across $t\in[T]$ and suppose $\max_{t,x}\,\big\{\frac{p_{t}^{\,u}(x)}{p_{t}^{\,r}(x)}\!\vee\!\frac{p_{t}^{\,r}(x)}{p_{t}^{\,u}(x)}\big\}=:M<\infty$. Let $C:=\max_{t,x:\,p_{t}^{\,u}(x)>0}\bigl[\log\frac{p_{t}^{\,r}(x)}{p_{t}^{\,u}(x)}\bigr]^{2}<\infty$. Then, for any ${\varepsilon}>0$, with probability at least $1-2\exp\bigl(-T{\varepsilon}^{2}/(2C)\bigr)$,
We start with the following three lemmata that are needed for the proof of Theorem B.1:
Lemma B.1 (Point-wise KL bound)
@@
Proof Fix $x\in V$ and set $y:=\dfrac{p(x)}{q(x)}\in[1,M]$. Inequality equation 4 is equivalent to
@@
Multiplying both sides by y1y-1 yields equation 4, which is precisely equation 4 after reinstating $y=p(x)/q(x)$. Therefore, equation 4 holds for every $x\in V$. This completes the proof.
Lemma B.2
@@
Proof Let $m=\tfrac{p+q}{2}$. Pinsker’s inequality gives $\|p-m\|_{1}^{2}\leq 2\,D_{\mathrm{KL}}(p\|m)$ and analogously for $q$. Hence
@@
Lemma B.3 (Exact TV scaling under mixture)
@@
Proof $p^{d}-p^{r}=(1-\alpha)(p^{u}-p^{r})$. Taking $\ell_{1}$-norms and dividing by $2$ yields the identity.
Now, we are ready to prove Theorem B.1:
Proof Define $Y_{t}:=\log\frac{p_{t}^{\,r}(U_{t})}{p_{t}^{\,u}(U_{t})}$, so that
@@
By Jensen’s inequality, $\frac{1}{T}\sum_{t}\sqrt{\operatorname{\mathrm{JSD}}(p_{t}^{\,d},p_{t}^{\,r})}\leq\sqrt{\frac{1}{T}\sum_{t}\operatorname{\mathrm{JSD}}(p_{t}^{\,d},p_{t}^{\,r})}.$ Combining the displays proves the claim with $I(X_{\operatorname{MarI}};Z)=\frac{1}{T}\sum_{t}\operatorname{\mathrm{JSD}}(p_{t}^{\,d},p_{t}^{\,r})$.
@@ Appendix C Appendix of Section 3
C.1 Theoretical error bound between position-wise vs. pooled MarI
Here, we provide the theoretical analysis of the error between MarI and pooled MarI. The following result shows that, under a mild assumption, the error of using the pooled MarI to estimate MarI is bounded by the sequence-wise density variance:
Theorem C.1 (Pooling error bound)
For each $t$, set $m_{t}:=\tfrac{1}{2}(p_{t}^{\,d}+p_{t}^{\,r})$ and $\bar{m}:=\tfrac{1}{2}(\bar{p}^{\,d}+\bar{p}^{\,r})$, where $\bar{p}^{\,d}:=\tfrac{1}{T}\sum_{t=1}^{T}p_{t}^{\,d}$ and $\bar{p}^{\,r}:=\tfrac{1}{T}\sum_{t=1}^{T}p_{t}^{\,r}$. Assume the uniform overlap condition
$\beta\ :=\ \min\Bigg\{$ $\inf_{\lambda\in[0,1]}\min_{t,x}\!\bigl[(1-\lambda)\,m_{t}(x)+\lambda\,\bar{m}(x)\bigr],$ (5)
$\inf_{\lambda\in[0,1]}\min_{t,x}\!\bigl[(1-\lambda)\,p_{t}^{\,d}(x)+\lambda\,\bar{p}^{\,d}(x)\bigr],\inf_{\lambda\in[0,1]}\min_{t,x}\!\bigl[(1-\lambda)\,p_{t}^{\,r}(x)+\lambda\,\bar{p}^{\,r}(x)\bigr]\Bigg\}\ >\ 0.$ (6)
Define the (averaged) $\ell_{2}$-deviation terms
$V_{d}\ :=\ \frac{1}{T}\sum_{t=1}^{T}\bigl\|p_{t}^{\,d}-\bar{p}^{\,d}\bigr\|_{2}^{2},\qquad V_{r}\ :=\ \frac{1}{T}\sum_{t=1}^{T}\bigl\|p_{t}^{\,r}-\bar{p}^{\,r}\bigr\|_{2}^{2}.$
Then
$0\ \leq\ I(X_{\operatorname{MarI}};Z)\;-\;I(\bar{X}_{\operatorname{MarI}};Z)\ \leq\ \frac{1}{4\,\beta}\,\Bigl(V_{d}+V_{r}\Bigr),$ (7)
where $I(X_{\operatorname{MarI}};Z)=\tfrac{1}{T}\sum_{t=1}^{T}\operatorname{\mathrm{JSD}}(p_{t}^{\,d},p_{t}^{\,r})$ and $I(\bar{X}_{\operatorname{MarI}};Z)=\operatorname{\mathrm{JSD}}(\bar{p}^{\,d},\bar{p}^{\,r})$.
Proof The lower bound $I(\bar{X}_{\operatorname{MarI}};Z)\leq I(X_{\operatorname{MarI}};Z)$ follows directly from the data-processing inequality. For the upper bound, write
$\frac{1}{T}\sum_{t=1}^{T}\operatorname{\mathrm{JSD}}(p_{t}^{\,d},p_{t}^{\,r})\;-\;\operatorname{\mathrm{JSD}}(\bar{p}^{\,d},\bar{p}^{\,r})$
$=$ $\frac{1}{T}\sum_{t=1}^{T}\Bigl[\,\underbrace{H(m_{t})-H(\bar{m})}_{(A_{t})}-\frac{1}{2}\underbrace{\bigl(H(p_{t}^{\,d})-H(\bar{p}^{\,d})\bigr)}_{(B_{t})}-\frac{1}{2}\underbrace{\bigl(H(p_{t}^{\,r})-H(\bar{p}^{\,r})\bigr)}_{(C_{t})}\Bigr].$
Let $\Delta_{m}:=m_{t}-\bar{m}$, $\Delta_{d}:=p_{t}^{\,d}-\bar{p}^{\,d}$, $\Delta_{r}:=p_{t}^{\,r}-\bar{p}^{\,r}$. Since $H$ is twice-differentiable, the second-order Taylor expansion around the pooled densities yields
$H(a)\;=\;H(a^{\prime})+\langle\nabla H(a^{\prime}),a-a^{\prime}\rangle+\tfrac{1}{2}(a-a^{\prime})^{\top}\nabla^{2}H(s)(a-a^{\prime}),$
for some $s$ on the line segment between $a^{\prime}$ and $a$. Note that $H$ is concave, $\nabla^{2}H(s)$ is negative semidefinite and diagonal with entries 1/s(x)-1/s(x). By the overlap assumption equation 5, every coordinate along the segments between $m_{t}$ and $\bar{m}$ and between $p_{t}^{\,s}$ and $\bar{p}^{\,s}$ ($s\in\{d,r\}$) is at least $\beta$, hence
$(a-a^{\prime})^{\top}\nabla^{2}H(s)(a-a^{\prime})\ \leq\ \frac{1}{\beta}\sum_{x\in V}\bigl(a(x)-a^{\prime}(x)\bigr)^{2}\ \leq\ \frac{1}{\beta}\,\|a-a^{\prime}\|_{2}^{2},$
and therefore the (negative) Taylor remainders satisfy
$H(a)-H(a^{\prime})-\langle\nabla H(a^{\prime}),a-a^{\prime}\rangle\ \geq\ -\,\frac{1}{2\beta}\,\|a-a^{\prime}\|_{2}^{2}.$ (8)
Applying equation 8 to the three entropy differences and averaging in $t$, the first-order terms vanish because $\frac{1}{T}\sum_{t}\Delta_{d}=0$, $\frac{1}{T}\sum_{t}\Delta_{r}=0$, and $\frac{1}{T}\sum_{t}\Delta_{m}=0$. Thus,
$\frac{1}{T}\sum_{t}\bigl[(A_{t})-(B_{t})/2-(C_{t})/2\bigr]\leq$ $\frac{1}{T}\sum_{t}\frac{1}{2}\Bigl(-\mathcal{R}_{d}(t)\Bigr)\;+\;\frac{1}{T}\sum_{t}\frac{1}{2}\Bigl(-\mathcal{R}_{r}(t)\Bigr)-\frac{1}{T}\sum_{t}\mathcal{R}_{m}(t)$
$\leq$ $-\frac{1}{2}\Big(\frac{1}{T}\sum_{t}\Bigl(\mathcal{R}_{d}(t)\Bigr)\;+\;\frac{1}{T}\sum_{t}\Bigl(\mathcal{R}_{r}(t)\Bigr)\Big)$
[…]
$=$ $\frac{1}{4\beta}\,(V_{d}+V_{r}).$
where $\mathcal{R}_{s}(t):=H(p_{t}^{s})-H(\bar{p}^{s})-\langle\nabla H(\bar{p}^{s}),p_{t}^{s}-\bar{p}^{s}\rangle\leq 0$ for $s\in\{d,r\}$ and $\mathcal{R}_{m}(t):=H(m_{t})-H(\bar{m})-\langle\nabla H(\bar{m}),m_{t}-\bar{m}\rangle\leq 0$.
C.2 Theoretical Guarantee provided by pooled MarI
In general, there is no non-trivial (distribution–free) theoretical guarantee for the sequence–level perplexity gaps can be stated solely in terms of the pooled MarI $I(\bar{X}_{\operatorname{MarI}};Z)=\operatorname{\mathrm{JSD}}(\bar{p}^{\,d},\bar{p}^{\,r})$ without any additional assumption, such as the bounded $\ell^{2}$-deviation in Theorem C.1 above.
Nonetheless, the pooled MarI can certify the following word–level forgetting (for particular tokens) forgetting:
Theorem C.2 (Word-level provable unlearning via pooled MarI)
Fix a token $w\in V$ and assume $\bar{p}^{\,u}(w)\wedge\bar{p}^{\,r}(w)=:\bar{\gamma}_{w}\in(0,1]$. Then
$\Big|\,\log\bar{p}^{\,u}(w)-\log\bar{p}^{\,r}(w)\,\Big|\;\leq\;\frac{2}{\bar{\gamma}_{w}\,(1-\alpha)}\;\sqrt{\,2\,I(\bar{X}_{\operatorname{MarI}};Z)\,}.$
Proof It follows from mean value theorem for $x\mapsto\log x$ on $\bar{\gamma}_{w},1]$ that
$|\log\bar{p}^{\,u}(w)-\log\bar{p}^{\,r}(w)|\leq|\bar{p}^{\,u}(w)-\bar{p}^{\,r}(w)|/\bar{\gamma}_{w}.$
Since $\bar{p}^{\,u}-\bar{p}^{\,r}=(\bar{p}^{\,d}-\bar{p}^{\,r})/(1-\alpha)$, we have
$|\bar{p}^{\,u}(w)-\bar{p}^{\,r}(w)|\;\leq\;\|\bar{p}^{\,u}-\bar{p}^{\,r}\|_{1}=\frac{1}{1-\alpha}\,\|\bar{p}^{\,d}-\bar{p}^{\,r}\|_{1}=\frac{2}{1-\alpha}\,\|\bar{p}^{\,d}-\bar{p}^{\,r}\|_{TV}.$
Finally, combining the above two inequalities, we have
$|\log\bar{p}^{\,u}(w)-\log\bar{p}^{\,r}(w)|\leq\frac{2}{\bar{\gamma}_{w}(1-\alpha)}\,\|\bar{p}^{\,d}-\bar{p}^{\,r}\|_{TV}\leq\frac{2}{\bar{\gamma}_{w}(1-\alpha)}\,\sqrt{2\,I(\bar{X}_{\operatorname{MarI}};Z)},$
where the last inequality follows from the fact that
$\|\bar{p}^{\,d}-\bar{p}^{\,r}\|_{TV}\leq\sqrt{2\,\operatorname{\mathrm{JSD}}(\bar{p}^{\,d},\bar{p}^{\,r})}=\sqrt{2\,I(\bar{X}_{\operatorname{MarI}};Z)}.$
C.3 Empirical error bound between position-wise vs. pooled MarI
@@ C.2 Theoretical Guarantee provided by pooled MarI
In our experiments, because the text is heterogeneous in both length and context, we use random mini-batches and the pooled estimator by default.
@@ Appendix D Appendix of Section 4
D.1 GPU, Computation and the Algorithm
[…]
Method Time per batch (s / batch) Peak memory (GB / GPU)
F-MarI 0.56 34.19
KL-GA 0.50 31.22
GD 0.45 27.15
DPO 0.57 36.35
GA 0.32 28.42
Table 4: Compute cost for different unlearning methods on GPT-2 Large. We report the average time per batch, and peak per-GPU memory usage.
During unlearning of the CP using Llama, the batch size was 3 for the unlearn set and 6 for the retain set. All experiments were conducted using 2 × NVIDIA GRID T4-16Q 16GB GPUs at maximum memory utilization.
Method Time per batch (s / batch) Peak memory (GB / GPU)
F-MarI 1.95 $\sim$16
KL-GA 1.93 $\sim$16
GD 1.50 $\sim$16
DPO 1.76 $\sim$16
GA 0.63 $\sim$16
D.2 Full Algorithm and Flow Chart
Here, we first provide both the full-detailed algorithm pseudo-code for Forgetting-MarI and the flowchart for readers who are more familiar with chart presentations.
Figure 8: Flow chart for Forgetting-MarI.
Algorithm 1 Forgetting-MarI.
Here, we note that, in practice, encoding often introduces padding tokens, and one should ignore those for downstream calculations: $X_{R}^{\text{flat}}\leftarrow\operatorname{flatten}\,(L_{R}[i:x_{i}\neq\text{pad}])\in\mathbb{R}^{N_{r}\times V}$; and do similarly for $X_{U}^{\text{flat}}$. This ensures that probabilities derived from $L_{R}^{\text{flat}}$ are not biased by padding positions.
D.3 Ablation Study for the GPT2-LG
In this ablation study, we deliberately curated the datasets to be maximally correlated, creating conditions where the distinction between forget and retain information is most difficult. This setup serves as a stress test for unlearning methods, as it requires the algorithm to selectively remove knowledge that is deeply intertwined with information that should be preserved.
Different $\gamma$ Figure 10 reports the training curves of all the compared full-parameter tuning UL methods using different regularization parameters $\gamma$. Due to the marginal information unlearning nature, Forgetting-MarI has the advantage that a wide range of parameter choice results in fast convergence around the unlearn baseline, indicating robust parameter-tuning. In comparison, it is clear that other methods demonstrate non-convergent (unstable) learning trajectory with extremely narrow parameter range to get close to the unlearn baseline, indicating extreme difficulty and effort in parameter-tuning.
Learning Rate Figure 9 tests a larger learning rate scenario among all methods.
These results collectively demonstrate that Forgetting-MarI provides more precise control over the unlearning process, maintaining the critical balance between forgetting targeted information and preserving general utility. The results highlights two notions of robustness:
Figure 9: Training curves for each method with varying choices of the regularization parameter $\gamma$ and lr=1e-4. Forgetting-MarI exhibits smooth monotone behavior, while the other methods show oscillation or utility collapse.
Figure 10: Training Curves for full-parameter UL methods with different $\gamma$ choices with lr=1=1e-5.
(1) Training robustness over epochs. Forgetting-MarI descends steadily to its optimum as seen in Fig. 9. GA and GD overshoot and bounce, KL–GA diverges after 5–6 epochs, DPO plateaus prematurely.
(2) Robustness against regularization tuning. Forgetting-MarI shows a monotone and smooth utility-unlearning trade-off when adjusting the regularization parameter. In comparison, GD and KL-GA display unstable oscillations with different choices of $\gamma$.
Both training and regularization robustness are necessary for a practical use of unlearning techniques. Practitioners do not have access to ground truth baselines and have limited time to select/determine the best parameter or training epoch to stop at, so stability is essential. Forgetting-MarI is the most stable technique during unlearning, making it the safest choice in practice.
D.4 Supplemental to Sec 4.2
Stability of continual unlearning Figure 11 below shows the unlearning trajectories of different methods on GPT-2 with Harry Porter and Llama 3.2-1B with Careless People. The unlearning performance of each method over the course of unlearning, where the curves for each method correspond to the experiment with the best performing regularization parameter for each method.
Figure 11: Next-token prediction accuracy during unlearning training across different methods. Horizontal dashed lines represent the “gold standard” unlearn baseline (model trained only on the retain dataset $\mathcal{D}_{r}$). (a) Top: Results for GPT-2 Large on Harry Potter, showing training dynamics across epochs (epoch 0 represents pre-unlearning performance). (b) Bottom: Results for Llama-3.2-1B on Careless People, with left and right columns for each method showing correlated and uncorrelated test settings, respectively.
Figure 12: MI Loss for GPT-2 Large on Harry Potter Dataset. Here, we observe nearly perfect correlation between the designed marginal information regularization loss and the unlearn accuracy, under different choice of regularization parameter.)
We note that there is the possibility that, in theory, one could find a perfect balance between gradient ascent and utility regularization, leading to a stable balance between unlearning and utility preservation, using one of the other methods. However, such a balance seems practically unattainable due to the unlearning instability over time and the lack of monotonicity in the choice of $\gamma$ for methods based on gradient ascent.
Figure 13: Training curves of F-MarI for the 10/90 split unlearning with the GPT2-LG model.
Additionally, we report the training curves for various $\gamma$ for F-MarI for the HP in Figure 13. We also plot the proposed MI loss across a few $\gamma$.
Finally, in Figure 12, we report an observation of nearly perfect correlation between the designed marginal information regularization loss and the unlearn accuracy, under different choice of regularization parameter.
D.5 Supplementary General Model Capacity Test Results